A curve in a road has a radius of 240 m. If the road surface is level and on a  

               rainy day the coefficient of static friction is 0.48, what is the maximum speed 

               at which the car can negotiate this curve? 

            

         


1
Expert's answer
2022-05-30T15:18:15-0400
"ma=F"

"\frac{mv^2}{R}=\mu m g"

"v=\sqrt{\mu g R}\\
=\sqrt{0.48*9.8*240}=33.6\:\rm m/s"